Suppose that Second Bank of Maine holds securities for $100 and its loans portfolio is valued at $200. Off-balance-sheet activities generated income for $20 this year. If this bank earns 5 percent on its securities and 10 percent on its loans, its interest income for this year is:

Also what is its non-interest income for this year?
Answer #1

Interest income for this year = ($100 * 0.05) + ($200 * 0.10) = 5 + 20 = $25

Its non-interest income for this year = $20